# Towards Fully Autonomous Research Powered by LLMs: Case Study on Simulations Zhihan Liu,a Yubo Chai,a Jianfeng Li,a

### v2

@@ Discussion
The experiments demonstrate that LLMs, with their robust coding ability 24, 32, 33, 34, can be directly employed in developing agents to autonomously complete complex research missions when aided by proper API programming. ASA-GPT-4o and ASA-GPT-4-turbo, notably, exhibits near-perfect completion rates for simpler missions, highlighting their potential. Multi-tier RP designs effectively handle more complex missions, though accurate information transmission remains challenging for such RPs. We emphasize that these simulation tasks, particularly those involving remote server execution, cannot be accomplished through interactions with web-based LLMs alone.
@@
ASA achieves fully autonomous research through continuous interaction with the LLM according to a specific logical design guided by the research plan (RP) provided by the researchers. We have identified several key practices in ASA logic design and RP design to enhance execution success rates.
  • Structure agent responses for efficient code and task extraction. Require the LLM to produce structured outputs, enabling ASA to accurately extract Python code, subtasks, and progress indicators. This improves code execution success rates and ensures smooth mission progression.
  • Incorporate agent-led code review and debugging. While the LLM is powerful in code generation, it may struggle with complex programs on the first attempt. By integrating self-review and debugging process, the LLM can generate correct and task-compliant code within several dialogue rounds.
@@
  • Provide detailed information for complex operations like remote server connections. Providing sufficient information in RP can improve mission success rates. In addition to necessary numeric/textual information, sometimes offering specific methods can guide the LLM to write correct programs more efficiently.
In summary, while the LLM has strong generative capabilities, its unpredictable outputs can complicate automation. ASA, based on Python’s AutoProg, though less intelligent, is reliable when logic is well-designed. By effectively structuring logic, ASA can oversee the LLM’s output, enhancing process reliability and achieving fully autonomous research.
Fig.  4: Recommended practices for fully autonomous research. By conducting iterative dialogues with the LLM adhering to a specific design logic, the success rate and efficiency of the automated process are enhanced. The diagrams on the left and right illustrate scenarios without and with the implementation of these strategies, respectively.
@@
Despite varying completion rates among tested ASAs, their errors and failure points share common characteristics, reflecting the inherent logical features of LLMs.
First, some ASAs tend to overlook parts of the RP’s requirements or “take shortcuts”, not fully following the RP’s instructions. For instance, they might skip plotting, fail to write a report, or ask humans to execute part of the task. These issues affect mission completeness and can disrupt the entire workflow. To improve autonomy in AI for science, we need to address these issues by continuously generating outputs describing the current mission status, periodically reviewing generated data, and confirming strict adherence to mission requirements. These improvements can be integrated through AutoProg modifications or by incorporating them into the "system" prompt input to LLMs or fine-tuning the LLMs.
Second, ASAs often underperform on technical aspects of tasks due to a lack of specialized domain knowledge. For example, when simulating a random walk, the algorithm needs to generate unit vectors uniformly distributed on a sphere to ensure isotropy. Although ASAs can often generate spatially random unit vectors, they fail to ensure uniform distribution on the sphere. In another example, for simulating self-avoiding walks, ASAs often incorrectly extend the random walk strategy by merely checking distances between points, resulting in high-energy state samples with large statistical errors. Correct sampling should use importance sampling to generate statistically significant conformations35. To avoid these issues in research, reliable solutions should be provided in the RPs, or relevant domain knowledge should be used for fine-tuning the LLMs to enhance their understanding.
@@
Each AutoProg involves a Python code debugging process. When bugs are detected, the LLM receives error messages and attempts to revise the code, iterating until it executes correctly. However, we’ve noted instances where the LLM struggles to generate functional code, resulting in mission failures. This typically occurs when the LLM becomes “stuck” after repeated unsuccessful debugging attempts, cycling through the same erroneous code without progress.
This phenomenon may be associated with the LLM’s propensity to depend heavily on past dialogue content. Research has documented instances where LLMs utilize context to deliver responses preferred by users, indicating an excessive reliance on previous information 36, 37, 38. While this capability allows LLMs to adapt responses based on short dialogue histories, it can also lead to persistent errors.
In our experiments, to prevent such deadlocks, we implemented a maximum debug attempt limit. Once this limit is reached, the AutoProg returns a “mission failed” message. To ensure the smooth execution of concrete research missions, it is crucial to avoid these debugging loops. One solution involves clearing parts of the memory after a certain number of debug attempts or encouraging the LLM to generate varied content in each iteration (Some LLMs include parameters to avoid generating duplicate content and to adjust the creativity of the models.).
@@
In our experiments, ASAs exhibited a significant level of global oversight, which allows them to track mission progress based on the dialogue history. They can retain details from earlier stages of missions even after focusing on lengthy sub-tasks. However, as discussed above, ASAs occasionally skip steps in prolonged missions or fail to deliver comprehensive solutions for sub-tasks, indicating a lack of sufficient local attention.
@@
In Automating Agent Coordination section, we designed a collaboration plan where a Main AI and some Subordinate AIs handled different aspects of the mission: the Main AI defined sub-task requirements and received reports, while Subordinate AIs focused on specific sub-tasks. This approach minimized redundant information in their dialogue history, enhancing the Main AI’s global awareness and the Subordinate AIs’ local attention. While the Main AI effectively completed the mission in RP 1, the Main-Sub AI model encountered challenges in RP 2 due to the failure of precise information transmission. One potential optimization is to involve human-designed RPs distributed among various Subordinate AIs for execution.
@@
This study demonstrates the capability of LLM powered research agent ASA to independently undertake sophisticated simulation research missions, using a well-recognized polymer physics simulation problem and a more challenging celestial-simulation problem as representative examples. However, there remains ample opportunity to explore ASA’s potential in tackling novel scientific challenges. Additionally, this study focuses on presenting an automated workflow, leaving room for further optimization of prompts and automation programs.
@@
Several inherent limitations of ASAs were identified: (1) ASAs may overlook mission requirements, necessitating methods to prioritize critical information; (2) general LLMs lack specialized domain knowledge, requiring supplementary information in RPs; (3) LLMs may persist in errors by relying too heavily on previous content, requiring memory management strategies in ASAs’ development; (4) balancing global awareness and local attention poses a challenge; (5) accurately transmitting specific information remains problematic.
Addressing these issues is crucial for enhancing the role of AI in scientific research. This requires a thorough analysis of these limitations and adjustments to research methodologies and LLM capabilities, leveraging their strengths while mitigating their weaknesses.
Conclusion
This study demonstrates the potential of LLM powered research agent ASA to autonomously conduct scientific research missions, using a polymer theory simulation problem and a more challenging asteroid-trajectory prediction problem as model problems. This process, encompassing simulation coding, remote execution, data analysis, and report generation, represents the first fully autonomous AI-directed simulation research: once the research plan is made, only a single command (e.g., ‘do prj.txt’) is required to initiate and complete the process. The notable success herein can be largely attributed to the sophisticated API programming. Agents such as ASA-GPT-4o demonstrate high completion rates for missions, a strong grasp of task requirements, and consistent alignment with overall objectives across extensive tasks. Notably, the performance exhibited in completing RP 4, entailing a hierarchy of nested tasks, surpassed expectations in a significant manner. These capabilities underscore the readiness of ASAs to independently manage long-term scientific research missions, offering a new perspective on AI applications in science as well as in chemical engineering.
@@ Methods
Based on the official documentation of various LLMs, we implemented the API automation programs (AutoProg) by configuring Python API calling methods and setting up the local environment. This setup included installing necessary Python packages like numpy, matplotlib, paramiko, and python-docx tailored for our simulation tasks. An AutoProg serves dual functions: 1) it engages in dialogue with the LLM via the API and preserves all historical dialogues, and 2) it extracts and executes Python programs generated by the LLM in response to prompts.
For instance, in the case of RP 1, an AutoProg initiates the automated research process by reading mission requirements saved in a txt file. It then calls the API to transmit the RP to the LLM, instructing it to generate a comprehensive Python program. The AutoProg extracts this program from the LLM’s response, sends it back to the LLM via the API along with entire dialogue history, and requests error checking and compliance verification against mission requirements. Subsequently, the AutoProg executes the validated Python program, capturing and storing output or errors. In case of errors, the AutoProg sends the prior dialogues and program errors back to the LLM, asking for modifications. This process repeats until the program runs correctly. If errors persist beyond a predefined threshold, the AutoProg halts with a “mission failed” output. Conversely, upon successful execution, it returns the program output to the LLM, prompting it to check the dialogues and files in the working directory to determine the task progress. If the task remains incomplete, the next Python program will be generated by the LLM to continue the mission, iterating through checking and debugging processes until completion. When the mission achieves full execution, the AutoProg concludes with a "mission complete" output.
AutoProg.py and RP.txt, containing predefined instructions, are stored in the local working directory. To initiate the automated research process for each experiment, users navigate to the directory via the command line (CMD) and execute commands such as “python AutoProg.py -s RP.txt”. Alternatively, AutoProg can be elevated to a command by creating a batch file (Windows) or shell script (Linux), allowing users to simply enter “AutoProg RP.txt” from any directory to execute it. A detailed video recording of the execution process is available for reference (SI-video), in which a batch file is used to facilitate the execution by running the command “do AI4S_prj1.txt”.
To evaluate LLM performance across RP 1-3, each ASA underwent 20 tests per RP, with all generated files and AutoProg outputs saved. Sample results are provided in the data (SI-data-1).
@@
The EWM, a well-established approach for determining the relative significance of criteria in decision-making 39, was utilized to assign weights based on data variability. This ensures criteria with higher variability, and thus more informational value, receive higher weights in the final evaluation.
TOPSIS, originally formulated by Hwang and Yoon 40, was subsequently applied. This method computes the geometric distance between each agent’s performance and both ideal and anti-ideal solutions, facilitating a comparative assessment of ASA efficacy.
@@
Initially, fulfillment data for all ASAs across various criteria were gathered over 20 trials, encapsulated in matrix X, wherein $x_{ij}$ signifies the frequency of criterion j’s attainment by agent i. Normalization was then performed on $x_{ij}$:
@@
yielding the normalized matrix $\mathbf{R}=[r_{ij}]$. Subsequently, criterion weights were established utilizing EWM. The relative proportion $p_{ij}$ of the criterion j:
@@
with $k=1/\ln(m)$, and $m$ denoting the total number of agents.
The weight $\omega_{j}$ of the criterion $j$:
@@
Then the geometric distances $S_{j}^{+}$ and $S_{j}^{-}$ from $v_{ij}$ to the positive and negative ideal solutions can be computed, respectively, as
@@
The relative closeness $C_{i}^{*}$ for each agent was calculated as:
@@
The following Supporting Information is available free of charge.
Partial results for gravitational simulation mission along with examples of reports written by ASAs (SI.pdf); partial AutoProgs and experimental result files for RP 1-3, main AI generating RP, and Multi-tier human RP (Some information has been obscured; readers can replace it with their own information and follow the instructions to run the AutoProgs) (SI-data-1.zip and SI-data-2.zip); a demo video demonstrating the operation process of the automatic research system (SI-video.mp4).
@@
Z.H.L. was the primary author of the paper. Y.B.C. was responsible for part of the data analysis and statistics. J.F.L. originally initiated and supervised the project.
@@ ACKNOWLEDGEMENTS
J.F.L. acknowledge supports from National Natural Science Foundation of China (Nos. 52394272, 22373022) and National Key Research and Development Program of China (No. 2023YFA0915300).
@@
  • Abolhasani and Kumacheva 2023 Abolhasani, M.; Kumacheva, E. The rise of self-driving labs in chemical and materials sciences. Nature Synthesis 2023, 2, 483–492.
  • Szymanski et al. 2023 Szymanski, N. J.; Rendy, B.; Fei, Y.; et al. An autonomous laboratory for the accelerated synthesis of novel materials. Nature 2023, 624, 86–91.
  • Volk et al. 2023 Volk, A. A.; Epps, R. W.; Yonemoto, D. T.; et al. AlphaFlow: autonomous discovery and optimization of multi-step chemistry using a self-driven fluidic lab guided by reinforcement learning. Nature Communications 2023, 14, 1403.

### v3

@@ Discussion
The experiments demonstrate that LLMs, with their robust coding ability,43, 51, 52, 53 can be directly employed in developing agents to autonomously complete complex research missions when aided by proper prompt engineering. ASA-GPT-4o and ASA-GPT-4-turbo, notably, exhibit near-perfect completion rates for simpler missions, highlighting their potential. Multitier RP designs effectively handle more complex missions, though accurate information transmission remains challenging for such RPs. We emphasize that these simulation tasks, particularly those involving remote server execution, cannot be accomplished through interactions with web-based LLMs alone.
@@
ASA achieves an autonomous research process guided by human-provided RPs through continuous interaction with the LLM according to a specific logical design. We identified several key practices in ASA logic design and RP design to enhance execution success rates.
  • Structure agent responses for efficient code and task extraction. Prompt the LLM to produce structured outputs, enabling ASA to accurately extract Python code, subtasks, and progress indicators. This improves code execution success rates and ensures smooth mission progression.
  • Incorporate agent-led code review and debugging. While the LLM is powerful in code generation, it may struggle with complex programs on the first attempt. By integrating self-review and debugging processes, ASA can generate correct and task-compliant code within several dialogue rounds.
@@
  • Provide detailed information for complex operations like remote server connections. Providing sufficient information in RP can improve mission success rates. In addition to necessary numeric/textual information, sometimes offering specific methods can guide ASA to write correct programs more efficiently.
In summary, while the LLM has strong generative capabilities, its unpredictable outputs can complicate automation. ASA, through prompt engineering to normalize LLM responses, although it restricts the LLM’s freedom, is reliable when logic is well-designed. By effectively structuring logic, ASA can oversee the LLM’s output, enhancing the smoothness and reliability of the automated simulation research process (Figure 4).
Fig.  4: Recommended practices for automated simulation research workflow. By conducting iterative dialogues with the LLM adhering to a specific design logic, the success rate and efficiency of the automated process are enhanced. The diagrams on the left and right illustrate scenarios without and with the implementation of these strategies, respectively.
@@
Despite varying completion rates among tested ASAs, their errors and failure points share common characteristics, reflecting the inherent logical features of the LLMs.
First, some ASAs tend to overlook parts of the RP’s requirements or “take shortcuts”, not fully following the RP’s instructions. For instance, they might skip plotting, fail to write a report, or ask humans to execute part of the task. These issues affect mission completeness and can disrupt the entire workflow. To improve the smoothness of the automated process, we need to address these issues by continuously generating outputs describing the current mission status, periodically reviewing generated data, and confirming strict adherence to mission requirements. These improvements can be integrated through prompt engineering modifications or by incorporating them into the “system” prompt input to LLMs or fine-tuning the LLMs.
Second, ASAs often underperform on technical aspects of tasks due to a lack of specialized domain knowledge. For example, when simulating a random walk, the algorithm needs to generate unit vectors uniformly distributed on a sphere to ensure isotropy. Although ASAs can often generate spatially random unit vectors, they fail to ensure uniform distribution on the sphere. In another example, for simulating self-avoiding walks, ASAs often incorrectly extend the random walk strategy by merely checking distances between points, resulting in high-energy state samples with large statistical errors. Correct sampling should use importance sampling to generate statistically significant conformations.54 To avoid these issues in research, reliable solutions should be provided in the RPs, or relevant domain knowledge should be used for fine-tuning the LLMs to enhance their understanding.
@@
The prompt engineering logic design of ASA includes a Python code debugging process. When bugs are detected, error messages are passed to the LLM, which is then prompted to revise the code. This process is iterative until the code executes correctly.
However, we have noted instances where ASA struggles to generate functional code, resulting in mission failures. This typically occurs when the LLM becomes “stuck” after repeated unsuccessful debugging attempts, cycling through the same erroneous code without progress.
This phenomenon may be associated with the LLM’s propensity to depend heavily on past dialogue content. Research has documented instances where LLMs utilize context to deliver responses preferred by users, indicating an excessive reliance on previous information.55, 56, 45 While this capability allows LLMs to adapt responses based on short dialogue histories, it can also lead to persistent errors.
In our experiments, to prevent such deadlocks, we implemented a maximum debug attempt limit. Once this limit is reached, ASA’s AutoProg returns a “mission failed” message. To ensure the smooth execution of concrete research missions, it is crucial to avoid these debugging loops. One solution involves clearing parts of the memory after a certain number of debug attempts or prompting the LLM to generate varied content in each iteration (Some LLMs include parameters to avoid generating duplicate content and to adjust the creativity of the models.).
@@
In our experiments, ASA exhibited a significant level of global oversight, which allows it to track mission progress based on the dialogue history. ASA can retain details from earlier stages of missions even after focusing on lengthy subtasks. However, as discussed above, ASA occasionally skips steps in prolonged missions or fails to deliver comprehensive solutions for subtasks, indicating a lack of sufficient local attention.
@@
In the Automating Agent Coordination through sub-RP Generation section, we designed a collaboration plan where a Manager AI and some Executor AIs handled different aspects of the mission: the Manager AI defined subtask requirements and received reports, while Executor AIs focused on specific subtasks. This approach minimized redundant information in different AIs’ dialogue history, enhancing the Manager AI’s global awareness and the Executor AIs’ local attention. While the Manager AI effectively completed the mission in RP 1, the Manager-Executor AI mode encountered challenges in RP 2 due to the failure of precise information transmission. One potential optimization involves human-designed RPs distributed among various Executor AIs for execution.
@@
This study demonstrates the capability of an autonomous simulation agent (ASA), constructed through LLM prompt engineering and an automated program, to automatically complete long-task workflows according to a Research Plan (RP) provided by humans. We used a well-recognized polymer physics simulation problem and a more challenging celestial simulation problem as representative examples.
@@
We tested different LLMs to construct ASAs, and while agents like ASA-GPT-4o, ASA-GPT-4-turbo, and ASA-Claude-3.5 demonstrated very high completion rates in the test tasks, we also identified several inherent limitations of the ASA method in handling long tasks: (1) Mission Requirement Overlooking. ASAs may overlook mission requirements, necessitating methods to prioritize critical information; (2) Lack of Specialized Domain Knowledge. General LLMs lack specialized domain knowledge, requiring supplementary information in RPs; (3) Error Persistence. LLMs may persist in errors by relying too heavily on previous content, requiring memory management strategies in ASAs’ development; (4) Balancing Global Awareness and Local Attention. Ensuring that ASA maintains a broad overview while focusing on specific details as needed poses a challenge; and (5) Information Transmission Accuracy. Securing that specific information is correctly interpreted and utilized remains problematic. These limitations highlight areas for further improvement and research to enhance the robustness and reliability of ASAs in handling complex, end-to-end simulation tasks.
Despite the current limitations of LLMs, which mean that the scientific conclusions drawn by ASA are not yet fully reliable, these results showcase the applicability of the ASA approach across various fields (Figures S4-S7) and its potential to assist in exploring new questions.
Further, we need to endow AI with more autonomous judgment and exploration capabilities, enabling it to summarize existing information and continuously generate new RPs for execution in more complex research scenarios, so as to tackle novel scientific challenges in the future.
Conclusions
This study demonstrates the potential of LLM-powered research agent ASA to autonomously conduct end-to-end scientific research missions under the guidance of an RP provided by humans. This comprehensive simulation research process, encompassing simulation coding, remote execution, data analysis, and report generation, represents an autonomous AI-conducted simulation research process. Once the research plan is made, only a single command (e.g., “do prj.txt”) is required to initiate and complete the process. The success herein can be largely attributed to prompt engineering and automated program design.
We tested ASAs equipped with different LLMs using both a well-known polymer chain simulation problem and a more challenging celestial simulation problem. Human-written RPs were provided, requiring the ASAs to execute according to their content, and the task completion of each agent was evaluated. Agents such as ASA-GPT-4o demonstrated high completion rates, a strong grasp of task requirements, and consistent alignment with the overall objectives across extensive tasks. Notably, the performance exhibited in completing RP 4, which entailed a hierarchy of nested tasks, surpassed expectations in a significant manner.
These capabilities underscore the readiness of ASA to conduct long-term research missions, offering a new perspective on AI applications in science, particularly in simulation studies, to enhance research efficiency.
@@ Methods
The ASA leverages an API automation program (AutoProg) to enable multiround text interactions with LLMs via API (Figure 1) and to execute the programs written by the LLM locally. AutoProg is designed with a logical prompt flow that guides the LLM to respond with specific content at different stages based on the conversation history and prompts, including prompting the LLM to break down the RP into subtasks, generate Python programs, debug the programs, and check the completion status of the tasks.
For instance, in the case of RP 1, ASA’s AutoProg initiates the automated simulation process by reading the RP.txt file written by a human. It then calls the API to transmit the content of the RP to the LLM, while prompting it to break the plan down into multiple subtasks and write Python code to complete the first subtask in its response. The AutoProg extracts the Python program from the LLM’s response, sends it back to the LLM via the API along with entire dialogue history, and requests error checking and compliance verification against mission requirements. Subsequently, AutoProg executes the validated Python program, capturing and storing output or errors. In the case of errors, the AutoProg sends the prior dialogues and program errors back to the LLM, asking for modifications. This process is repeated until the program runs correctly. If errors persist beyond a predefined threshold, the AutoProg halts with a “mission failed” output. Conversely, upon successful execution, it returns the program output to the LLM, prompting it to check the dialogues and files in the working directory to determine the task progress. If the task remains incomplete, it prompts the LLM to generate the next Python program to continue the mission, iterating through checking and debugging processes until completion. When the mission achieves full execution, the AutoProg concludes with a “mission complete” output.
AutoProg.py and RP.txt are stored in the local working directory. To initiate the automated simulation process using ASA for each experiment, users navigate to the directory via the command line (CMD) and execute commands such as “python AutoProg.py -s RP.txt”. Alternatively, AutoProg can be elevated to a command by creating a batch file (Windows) or shell script (Linux), allowing users to simply enter “command RP.txt” from any directory to execute it. A detailed video recording of the execution process is available for reference (SI-Video), in which a batch file is used to facilitate execution by running the command “do AI4S_prj1.txt”.
To evaluate the ASA performance powered by different LLMs across RP 1-3, each ASA underwent 20 tests per RP, with all generated files and outputs saved. Sample results are provided in the data (SI-data-1).
The ASA we developed is a general system (minor adjustments may be required in AutoProg for different LLM APIs) that requires only a change in the original RP to be applicable to any simulation problem involving Python programming. In addition to the RPs designed for polymer conformation simulations, we have also crafted RPs for solar system planetary orbit simulations and asteroid trajectory simulations (Figure 1C and SI).
@@
The EWM, a well-established approach for determining the relative significance of criteria in decision-making, 57 was utilized to assign weights based on data variability. This ensures criteria with higher variability, and thus more informational value, receive higher weights in the final evaluation.
TOPSIS, originally formulated by Hwang and Yoon,58 was subsequently applied. This method computes the geometric distance between each agent’s performance and both ideal and anti-ideal solutions, facilitating a comparative assessment of ASA efficacy.
@@
Initially, fulfillment data for all ASAs across various criteria were gathered over 20 trials, encapsulated in matrix X, wherein $x_{ij}$ signifies the frequency of criterion j’s attainment by agent i. Normalization was then performed on $x_{ij}$
@@
yielding the normalized matrix $\mathbf{R}=[r_{ij}]$. Subsequently, criterion weights were established utilizing EWM. The relative proportion $p_{ij}$ of criterion j
@@
with $k=1/\ln(m)$ and $m$ denoting the total number of agents.
The weight $\omega_{j}$ of the criterion $j$
@@
Then the geometric distances $S_{j}^{+}$ and $S_{j}^{-}$ from $v_{ij}$ to the positive and negative ideal solutions can be computed, respectively, as follows
@@
The relative closeness $C_{i}^{*}$ for each agent was calculated as follows
@@
Data Availability Statement
The source code and example results of ASA can be found at https://github.com/zokaraa/autonomous_simulation_agent.
@@
The following Supporting Information is available free of charge at https://pubs.acs.org/doi/10.1021/acs.jcim.4c01653.
  • Supporting diagram for Manager-Executor AI mode and Multitier AI mode; partial results for gravitational simulation mission; and results for AI-designed RP along with examples of reports written by ASAs (PDF)
  • A demo video demonstrating the operation process of the automatic research system (MP4)
  • ASA’s AutoProg examples and experimental result files for RP 1-3 (ZIP)
  • Manager-Executor AI mode, Multitier AI mode, and AI-designed RP (ZIP)
@@
Z.L. was the primary author of the paper. Y.C. was responsible for part of the data analysis and statistics. J.L. originally initiated and supervised the project.
@@ ACKNOWLEDGEMENTS
J.F.L. acknowledge supports from National Natural Science Foundation of China (Nos. 52394272, 22373022) , National Key Research and Development Program of China (No. 2023YFA0915300) and Shanghai Science and Technology Innovation Action Plan (No. 24JD1400700).
@@
  • Hu et al. 2024 Hu, J.; Zhou, L.; Jiang, J. Efficient Machine Learning Force Field for Large-Scale Molecular Simulations of Organic Systems. CCS Chemistry 2024, 1–15, Just Published.
  • Xu and Jiang 2024 Xu, L.; Jiang, J. Synergistic Integration of Physical Embedding and Machine Learning Enabling Precise and Reliable Force Field. Journal of Chemical Theory and Computation 2024, 20, 7785–7795.
  • 6 Li, J.; Zhang, H.; Chen, J. Z. Y. Structural Prediction and Inverse Design by a Strongly Correlated Neural Network. Physical Review Letters 123, 108002.
  • Liu et al. 2023 Liu, Z.; Liu, Y.-X.; Yang, Y.; Li, J. Template Design for Complex Block Copolymer Patterns Using a Machine Learning Method. ACS Applied Materials & Interfaces 2023, 15, 31049–31056.
  • Qiu and Sun 2024 Qiu, H.; Sun, Z. On-demand reverse design of polymers with PolyTAO. npj Computational Materials 2024, 10, 273.
  • Qiu et al. 2024 Qiu, H.; Wang, J.; Qiu, X.; Dai, X.; Sun, Z.-Y. Heat-Resistant Polymer Discovery by Utilizing Interpretable Graph Neural Network with Small Data. Macromolecules 2024, 57, 3515–3528.
  • Xu et al. 2022 Xu, F. F.; Alon, U.; Neubig, G.; Hellendoorn, V. J. A systematic evaluation of large language models of code. Proceedings of the 6th ACM SIGPLAN International Symposium on Machine Programming. New York, NY, USA, 2022; pp 1–10.
  • Hou et al. 2024 Hou, X.; Zhao, Y.; Liu, Y.; Yang, Z.; Wang, K.; Li, L.; Luo, X.; Lo, D.; Grundy, J.; Wang, H. Large Language Models for Software Engineering: A Systematic Literature Review. ACM Transactions on Software Engineering and Methodology 2024,
@@
  • Guo et al. 2024 Guo, E.; Gupta, M.; Deng, J.; Park, Y.-J.; Paget, M.; Naugler, C. Automated Paper Screening for Clinical Reviews Using Large Language Models: Data Analysis Study. Journal of Medical Internet Research 2024, 26, e48996.
  • Bolaños et al. 2024 Bolaños, F.; Salatino, A.; Osborne, F.; others Artificial intelligence for literature reviews: opportunities and challenges. Artificial Intelligence Review 2024, 57, 259.
  • Lozano et al. 2024 Lozano, A.; Fleming, S. L.; Chiang, C.-C.; Shah, N. Biocomputing 2024; 2024; pp 8–23.
  • Dagdelen et al. 2024 Dagdelen, J.; Dunn, A.; Lee, S.; others Structured information extraction from scientific text with large language models. Nature Communications 2024, 15, 1418.
  • Fink et al. 2023 Fink, M. A.; Bischoff, A.; Fink, C. A.; Moll, M.; Kroschke, J.; Dulz, L.; Heußel, C. P.; Kauczor, H.-U.; Weber, T. F. Potential of ChatGPT and GPT-4 for Data Mining of Free-Text CT Reports on Lung Cancer. Radiology 2023, 308, e231362.
  • Ma et al. 2023 Ma, P.; Ding, R.; Wang, S.; Han, S.; Zhang, D. InsightPilot: An LLM-Empowered Automated Data Exploration System. Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing: System Demonstrations. Singapore, 2023; pp 346–352.
  • Peng et al. 2024 Peng, D.; Zheng, L.; Liu, D.; others Large-language models facilitate discovery of the molecular signatures regulating sleep and activity. Nature Communications 2024, 15, 3685.
  • Li et al. 2024 Li, T.; Shetty, S.; Kamath, A.; others CancerGPT for few shot drug pair synergy prediction using large pretrained language models. npj Digital Medicine 2024, 7, 40.
  • Subba et al. 2024 Subba, B.; Toufiq, M.; Omi, F.; others Human-augmented large language model-driven selection of glutathione peroxidase 4 as a candidate blood transcriptional biomarker for circulating erythroid cells. Scientific Reports 2024, 14, 23225.
  • Liu et al. 2023 Liu, F.; Zhu, T.; Wu, X.; others A medical multimodal large language model for future pandemics. npj Digital Medicine 2023, 6, 226.
  • Madani et al. 2023 Madani, A.; Krause, B.; Greene, E. R.; others Large language models generate functional protein sequences across diverse families. Nature Biotechnology 2023, 41, 1099–1106.
@@
  • Yao et al. 2023 Yao, S.; Yu, D.; Zhao, J.; Shafran, I.; Griffiths, T.; Cao, Y.; Narasimhan, K. Tree of Thoughts: Deliberate Problem Solving with Large Language Models. Advances in Neural Information Processing Systems. 2023; pp 11809–11822.
  • Kojima et al. 2022 Kojima, T.; Gu, S. S.; Reid, M.; Matsuo, Y.; Iwasawa, Y. Large Language Models are Zero-Shot Reasoners. Advances in Neural Information Processing Systems. 2022; pp 22199–22213.
@@
For Table of Contents Use Only.
By submitting a research plan file using the command do prj1.txt, ASA will automatically execute all subsequent simulation tasks without any human intervention.